Martinez owns an asset that cost $87,000 with accumulated depreciation of $40,000. The company sells the equipment for cash of $42,000. At the time of sale, the company should record:

A gain on sale of $5,000.

A gain on sale of $2,000.

A loss on sale of $2,000.

A loss on sale of $45,000.

A loss on sale of $5,000.
Answer #1

Book value as on date of sale=Cost-Accumulated depreciation

=87,000-40,000=$47,000

Hence since sale proceeds is lesser than book value as on date of sale;

Hence loss =(47,000-42,000)=$5000

Hence the correct option is:

A loss on sale of $5,000.